### v3

@@ Abstract
TikTok is now a massive platform, and has a deep impact on global events. But for all the preliminary studies done on it, there are still issues with determining fundamental characteristics of the platform. We develop a method to extract a representative sample from a specific time range on TikTok, and use it to collect >99% of posts from a full hour on the platform, alongside a dataset of >99% of posts from a single minute from each hour of a day. Through this, we obtain post metadata, video media data, and comments from a close to complete slice of TikTok. Using this dataset, we report the critical statistics of the platform, notably estimating a total of 269 million posts produced on the day we looked at on TikTok, that 18% of videos on the platform feature children, and that $\sim$176 years of videos are uploaded every day.
keywords:
tiktok, sample, social-media, vertical-slice
@@ 1 Introduction
TikTok has reshaped global culture, marketing, and entertainment through its highly engaging, algorithm-driven platform that promotes short-form video content. Due to its immense reach [1], especially among younger voters [2, 3], and its ability to amplify content quickly through its algorithm, politicians, advocacy groups, and individuals use TikTok to share political messages, mobilize voters, and raise awareness about issues [4, 5].
However, concerns have been raised about the spread of misinformation [6], the platform’s potential for algorithmic bias [7], the lack of transparency in how political content is managed [8], and child safety on the platform [9]. Additionally, TikTok’s ownership and data practices have sparked debates over national security and foreign influence, particularly regarding its potential use in shaping public opinion during elections [10, 11, 12]. At the same time, little is known about the platform’s inner working and its global activity level. If we want to gain more insight into how TikTok feeds into these issues, we need strong quantitative insight into the platform.
Data collection methods on TikTok are limited. Search mechanisms on the site are limited to hashtag based search which stops at 1000 returns, fuzzy search, and user- and song-based search [13]. All of these methods limit the extent to which a true complete slice of the platform can be examined. While TikTok offers a Researcher API to selected scholars [14], it has been established recently that this API shows significant deviations in terms of metadata and availability of videos when compared to the TikTok website [15]. Additional limitations of the API include restrictive user agreements [16] and the inability to return videos originating from Canada  [^1] https://developers.tiktok.com/doc/research-api-codebook/. We must obtain stronger sampling methods if we are to gain contextualized insight into a platform of TikTok’s magnitude [17]. Without representative samples, researchers cannot properly contextualize or validate findings about behaviour on TikTok.
To address this need, we developed a novel method for collecting >99% uploaded to TikTok within a specific timespan, enabling a collection of a close-to-complete slice of TikTok for the first time.
We did this by reverse engineering how TikTok creates unique identifiers (IDs) for its videos, in order to create an ID generator that can allow us to search for all possible TikTok videos, which allows us to bypass the black box search elements normally available. Previous work using this method has allowed a greater understanding of YouTube [18].
We used the representative slice of TikTok that we collected to examine a number of fundamental phenomena on TikTok. Considering prevailing research themes around TikTok and other social media platforms, we provide insight on the following points:
@@
  • Distributions of post content on TikTok: We present data on country post origin, engagement statistics, and local posting time from our representative dataset. This allows us to understand global usage and consumption patterns, critical when considering the disparity between global populations and the TikTok user base [20].
@@
As part of this work, we also release a collection of fundamental statistics around these phenomena. We expect these will be useful to a wide-array of future studies: both for the studies themselves and to calibrate their findings against the broader backdrop of TikTok behavior.
In collecting the posts and media content for this study, we have violated TikTok’s Terms of Service  [^2] https://www.tiktok.com/legal/page/us/terms-of-service/en, specifically the following section: ‘use automated scripts to collect information from or otherwise interact with the Services;’ However, we strongly believe that our work is critical to better understand a platform that is used by billions of users worldwide. Greater platform transparency will lead to higher quality and more trustworthy research, which in turn, will help the community and TikTok work together to help the platform flourish.
@@ 2 Results
We developed a method that, to the best of our knowledge, allows us to collect >99% of the post content on TikTok across the world. We used that method to collect every post made between 5-6pm UTC on the 10th of April, 2024 (an arbitrary, mundane date, to obtain baseline data from the platform), and all posts from the 42nd minute of each hour from 10pm UTC on the 9th of April, till 10pm on the 10th of April, to obtain an across-day representative data sample. After collection, we collected the video media content and comments of the posts where available. Here, we use that data to extract what we consider to be the critical statistics of note on TikTok, that will let the research community contextualize findings they report on TikTok.
2.1 Statistics
We first present the high level statistics of our two datasets, enabling comparison with other dataset sizes.
Metric 1 Hour 1 Minute Each Hour
Number of Number of Videos 5,157,488 1,879,335
Number of Unique Users 3,715,672 1,760,335
Average Comment Count 4.92 4.18
Average Like Count 177.9 145.8
Average Share Count 14.1 9.14
Average View Count 2,533.2 2243.9
Average Video Duration (seconds) 20.69 20.42
Table 1: Statistics from the two datasets we collect for this work. Note that the two datasets overlap at the $42^{nd}$ minute of the hour we looked at in this work, so total counts are lower than the sum/mean of the two counts.
2.1.1 Estimating overall post volume
The nature of our dataset gives us first insight into the overall size of TikTok, a critical piece of knowledge for any social media platform [17].
For the observed hour on TikTok, we collected $5,157,488$ posts. By looking at error messages for other requested videos, we can determine that there were originally $12,044,235$ posts created in this hour (with many deleted, made private, or otherwise made inaccessible since post time, see Section 2.4), with an average of $3,345.6$ videos per second (minimum-maximum $2,773-9,042$, see Fig. 1(a)). The aggregated numbers per minute range from $189,620$ to $256,743$ (avg. $200,737.3$) videos per minute (Fig. 1(b)). Here, because we want to consider all posts that were available in the original time-span, included no longer available posts, we use the create time that can be extracted from the post ID, see Section 4.1.
In order to estimate the videos per hour for an entire day, we collected the videos of the $42^{nd}$ minute of every hour. For our collected hour, the $42^{nd}$ minute is slightly less active ($\sim 97.3\%$) than the average of the other $59$ minutes. Consequently, we used this correction factor to calculate the estimated number of videos per hour. The result can be seen in Fig. 1(c) and shows an interesting two-peaked temporal pattern. These estimated hours aggregate to 269.3 million video posts over a 24-hour time period.
@@ 2.1.1 Estimating overall post volume
We note that the time distribution of videos per hour looks markedly different from Corso et al. [14]. We see that there are 28.5% more videos in the first minute than in the average of the other minutes, implying scheduled posts.
Of videos that are still available when we collected data, and therefore that we have metadata for, we see that there is a 14% higher posting volume at the 0th second of each minute, implying scheduled posts, or bot activity. View counts for videos posted at the zeroth second get significantly more views (p-value: 7e-5), with an average of 3,899 views at the zeroth second vs. 2,502 views for all other seconds. Videos at the 0th second more frequently use hashtags implying the desire to reach a broader audience, such as fyp, implying that these posters are using post scheduling and hashtag use to attempt to reach a broader audience.
@@ 2.1.2 Engagement distributions
Figure 2: Histograms of various statistics in the video metadata, across the 1 hour dataset. We speculate that the peaks we see at round numbers on due to interval view count tracking optimization. However, they could also be evidence of inauthentic engagement.
Figure 2 shows the distribution of video statistics from the 1 hour dataset we collected. While comment count and share count follow a typical exponential distribution as previously seen [14], view count and like count show a mode at $\sim$100 views and $\sim$50 likes respectively, differing from the distributions reported by Corso et al. [14]. This may be due to the large discrepancies between the Research API and the data returned to the platform interface demonstrated by Pearson et al. [15], highlighting the need to double check Research API data with main platform data.
2.1.3 Understanding global usage
Figure 3: Choropleth of video posting per capita for all countries, determined via the locationCreated tag available on the video post metadata. We normalize by the population of each country to show posts per capita, which specifically accounts for population rather than citizens [22]. The counts are corrected for geographical regions that we estimate that we underrepresent.
Figure 4: Regression of country population against video count in this dataset, with the 6 countries furthest from the trend line labelled, as determined by highest and lowest residuals from the linear regression. Red points correspond to the country name closest to them.
We then looked at global usage of TikTok. Research on TikTok frequently looks at country-specific slices (e.g., [23]), so understanding actual country-level usage is critical for context [20]. We infer post-country-creation from the locationCreated tag that refers to the video’s origin country  [^3] https://developers.tiktok.com/doc/research-api-specs-query-videos. While this indicator of country origin is likely confounded by VPN usage and other factors, it represents our current best method for assessing a post’s country of origin. As detailed later in Section 4.1, our method has slight variation in regional representativeness, so we adjust counts based on known statistics from our method. We explain precisely how we do this in Section 4.3.
@@ 2.1.3 Understanding global usage
Figure 4 shows a regression plot of country population against video count in our dataset. We see a strong fit to the data, evidence that we have successfully collected a globally representative dataset. We see that China, India, and Cuba, have the least videos in our dataset per capita, based on the residual of a fitted linear regression. We note that while TikTok was previously extremely popular in India [14], there are now very few videos originating from India (as deemed by TikTok), likely due to the nationwide censorship policy on Chinese applications there [24]. Of course, there could still be videos produced from India via virtual private networks (VPNs), which would be tagged with a different origin country by TikTok. TikTok is not officially available in China [25], and Cuba has heavy internet restrictions [26]. Monaco, Qatar, and the UAE, have the most videos in our dataset per capita.
Using the locationCreated attribute, we can also infer the distribution of when TikTok posts are created in the local time, see Figure 5. In the case of countries spanning multiple timezones, we use the average hour from each timezone associated with the country from the Olson timezone database  [^5] https://github.com/stub42/pytz/blob/master/tz/zone.tab. Inaccuracies are inevitable due to lack of sub-country location information, and countries spanning multiple timezones, but the general trends should be preserved nonetheless.
Figure 5: Distribution of local time video posting frequency, found by looking at video post time combined with the locationCreated tag.
@@ 2.2 Video Media
Figure 6: View of the topics covered by the video posts, with each small dot representing a video, and distance between dots representing video content similarity. Common topics of videos are colored and labelled with a description describing the topic.
@@
The topic analysis completed here is not meant to be comprehensive. But by releasing the unsupervised topics discovered, we provide researchers an entrypoint to contemplating the study and engagement of particular domains within TikTok.
@@ 2.2.1 Presence of Children
The heavy use of TikTok by children and teenagers has caused concern for safety on the platform [9]. TikTok reports removing 445k accounts for child sexual exploitation and abuse  [^6] https://www.tiktok.com/transparency/en-us/child-exploitation over the first 6 months of 2024. But the lack of a representative sample of TikTok has meant so far we haven’t been able to estimate the true magnitude of child activity or the particular forms of child activity that are disproportionate on the platform. We know there is a high prevalence of children on the platform, with 50% of 5-7 year olds, and 80% of 12-15 year olds in the UK using the platform [21], but we don’t know what percentage of content on the platform features children. We aim to address that here.
@@
We trained a classifier to detect if a video contains a child, with a mean F1 of 0.86 from 5-fold cross-validation on our train/testset (precision 0.88, recall 0.81). We provide data, model selection, and training details in Section 4.4.2. We used this classifier to find that 18.13% of the videos in our 24 hour dataset contain children (95% confidence interval: [18.06%, 18.21%] from bootstrapped confidence intervals). This is a strikingly high statistic, and underscores how deeply embedded children are in platform activity. To the best of our knowledge, no comparable large-scale studies exist for TikTok or other platforms that would allow direct benchmarking or comparison of this figure.
@@
Figure 7: Choropleth of share of videos from our 1 minute per hour dataset containing one or more children, divided by the share of the population under the age of 14 [22] (to avoid showing simply a map of countries with a youthful population), plotted by country. Solid areas are statistically significant, hatched areas are not significant as determined by a binomial test (p=0.05).
@@
Figure 8: Scatter plot of countries, for each country showing the share of the population under the age of 14 against the share of videos from that country containing one or more children, as determined by our video classifier. We have removed countries where there was insufficient data to make a conclusive estimate of the share of videos containing children, based on a binomial test (p=0.05). We label the top 2 and bottom 2 countries with the highest residual from the linear regression, effectively showing countries with the most overrepresented and underrepresented children in TikTok posts, respectively. Red points correspond to the country name closest to them.
@@
We observe variability in the fraction of videos containing children by country. Figure 7 shows a choropleth of share of videos containing one or more children per country, adjusted by the share of the population under the age of 14  [^7] https://population.un.org/wpp/ (to avoid showing simply a map of countries with a youthful population). We removed countries where there were insufficient post counts to make a conclusive estimate of the share of videos containing children, based on a binomial test (p=0.05). This effectively shows countries where there is a culture of posting videos containing children. Figure 8 shows a scatter plot of countries comparing the share of the population under 14, and the share of videos containing one or more children - a linear trend is evident. We label the top 2 and bottom 2 countries with the highest residual from the linear regression, effectively showing countries with the most overrepresented and underrepresented children in TikTok posts, respectively. The variation we observe may reflect broader differences in cultural norms around public sharing of children’s images, the popularity of family-centered content, differences in platform access and usage patterns, and differences in the implementation of child protection policies across countries.

### v4

@@ Abstract
TikTok is now a massive platform, and has a deep impact on global events. Despite preliminary studies, issues remain in determining fundamental characteristics of the platform. We develop a method to extract a representative sample of >99% of posts from a given time range on TikTok, and use it to collect all posts from a full hour on the platform, alongside all posts from a single minute from each hour of a day. Through this, we obtain post metadata, video media, and comments from a close-to-complete slice of TikTok, and report the critical statistics of the platform. Notably, we estimate a total of 269 million posts produced on the day we looked at, that 18% of videos on the platform feature children, and that at least 0.5% of posts contain artificial intelligence-generated content.
@@ 1 Introduction
TikTok has reshaped global culture through its algorithm-driven platform that promotes short-form video content. Due to its immense reach [71], especially among younger voters [89, 96], and its ability to amplify content quickly through its algorithm, politicians, advocacy groups, and individuals use TikTok to share political messages, mobilize voters, and raise awareness about issues [98, 101].
However, concerns have been raised about the spread of misinformation [65], algorithmic bias [92], lack of transparency in how political content is managed [86], and child safety on the platform [119]. Additionally, TikTok’s ownership and data practices have sparked debates over national security and foreign influence, particularly regarding its potential use in shaping public opinion during elections [76, 120, 82]. At the same time, little is known about the platform’s inner workings and its global activity level. To gain insight here, we need good data from the platform.
Data collection methods on TikTok are limited [121], and do not allow for representative sampling. Search mechanisms on the site are limited to hashtag based search which stops at 1000 returns, fuzzy search, and user- and song-based search [121]. All of these methods limit the extent to which a true complete slice of the platform can be examined. While TikTok offers a Researcher API to selected scholars [77], it has been established recently that this API shows significant deviations in terms of metadata and availability of posts when compared to the TikTok website [108]. Additional limitations of the API include restrictive user agreements [66] and the inability to return posts originating from Canada  [^1] https://developers.tiktok.com/doc/research-api-codebook/. We must obtain stronger sampling methods if we are to gain contextualized insight into a platform of TikTok’s magnitude [110]. Without representative samples, researchers cannot properly contextualize or validate findings about behaviour on TikTok.
To address this, we developed a novel method for collecting >99% of posts uploaded within a timespan, enabling the collection of a close-to-complete slice of TikTok for the first time. We did this by determining how TikTok creates IDs for its posts, in order to create an ID generator that can allow us to search for all possible posts, which allows us to bypass the insufficient available search elements, as done previously on YouTube [126]. We expect this method and extensions of it to be invaluable to research efforts to understand the scope, dynamics, and impact of TikTok.
We used this method to collect near-complete samples of TikTok. Considering prevailing research themes around TikTok and other social media platforms, we provide insight on the following points:
@@
  • Distributions of post content on TikTok: We present data on country post origin, engagement statistics, and artificial-intelligence generated content from our representative dataset. This allows us to understand global usage and consumption patterns, critical when considering the disparity between global populations and the TikTok user base [103].
@@
In addition to releasing the code needed to reproduce this work, we also release fundamental statistics regarding these phenomena. We expect these will be useful to a wide-array of future studies: both for the studies themselves and to calibrate their findings against the broader backdrop of TikTok behavior. We invite researchers to approach us for specific data requests.
@@ 2 Results
We developed a method to collect >99% of TikTok’s global post content, and used it to collect posts made between 5-6pm UTC on the 10th of April, 2024 (an arbitrary, mundane date, to obtain baseline data), and posts from the $42^{nd}$ minute of each hour from 10pm UTC on the $9^{th}$ of April, till 10pm on the $10^{th}$ of April, to obtain an across-day representative sample. We first present the high-level statistics of our two datasets in Table 1.
Metric
1 Hour
1 Minute
Each Hour
No. of Videos 5,157,488 1,879,335
No. of Unique Users 3,715,672 1,760,335
Mean Comment Count 4.92 4.18
Mean Like Count 177.9 145.8
Mean Share Count 14.1 9.14
Mean View Count 2,533.2 2243.9
Mean Video Duration (s) 20.69 20.42
Table 1: Statistics from the two datasets we collected. The two datasets overlap at the $42^{nd}$ minute of the hour we looked at in this work, so total counts are lower than the sum of the two counts.
2.1 Overall post volume
The nature of our dataset offers us the most comprehensive assessment of the size of TikTok to date. For the observed hour, we collected 5 million posts. By looking at error messages for other requested posts, we determined that there were originally 12 million posts created in this hour (with many made inaccessible since post time, see Sec. 2.6), with a mean of 3.3k posts per second (min.-max. 2.8k9.0k2.8k-9.0k, see Fig. 1(a)). The aggregated numbers per minute range from $189,620$ to $256,743$ (mean $200,737.3$) posts per minute (Fig. 1(b)). Here, because we want to consider all posts that were available in the original time-span, included no longer available posts, we use the create time that can be extracted from the post ID, see Section 4.1.
In order to estimate the posts per hour for an entire day, we collected the posts of the $42^{nd}$ minute of every hour. For our collected hour, the $42^{nd}$ minute is slightly less active ($\sim 97.3\%$) than the mean of the other $59$ minutes. Consequently, we used this correction factor to calculate the estimated number of posts per hour. The result can be seen in Fig. 1(c) and shows an interesting two-peaked temporal pattern. These estimated hours aggregate to 269.3 million posts over a 24-hour time period.
@@ 2.1 Overall post volume
We note that the time distribution of posts per hour looks markedly different from [77]. We see that there are 28.5% more posts in the first minute than in the mean of the other minutes, implying scheduled posts.
Of posts that are still available when we collected data, and therefore that we have metadata for, we see that there is a 14% higher posting volume at the 0th second of each minute, implying scheduled posts, or bot activity. View counts for posts posted at the zeroth second get significantly more views (p-value: 7e-5), with an mean of 3,899 views at the zeroth second vs. 2,502 views for all other seconds. Videos at the 0th second more frequently use hashtags implying the desire to reach a broader audience, such as fyp, implying that these posters are using post scheduling and hashtag use to attempt to reach a broader audience.
@@ 2.2 Engagement
Figure 2: Histograms of various statistics in the post metadata, across the 1 hour dataset. We speculate that the peaks we see at round numbers on due to interval view count tracking optimization. However, they could also be evidence of inauthentic engagement.
Figure 2 shows the distribution of post statistics from the 1 hour dataset we collected. While comment count and share count follow a typical exponential distribution as previously seen [77], view count and like count show a mode at $\sim$100 views and $\sim$50 likes respectively, differing from the distributions reported by [77]. This may be due to the large discrepancies between the Research API and the data returned to the platform interface demonstrated by [108], highlighting the need to double-check Research API data with main platform data.
2.3 Global usage
Figure 3: Choropleth of posting per capita for all countries, determined via the locationCreated tag available on the post metadata. We normalize by the population of each country to show posts per capita, which specifically accounts for population rather than citizens [106]. The counts are corrected for geographical regions that we estimate that we under-represent. Areas in grey are countries where we had less than 50 posts available.
Figure 4: Regression of country population against post count in this dataset, with the 6 countries furthest from the trend line labelled, as determined by highest and lowest residuals from the linear regression. Red points correspond to the country name closest to them.
We then looked at global platform usage, inferring post origin country from the locationCreated attribute in the post metadata [^2] https://developers.tiktok.com/doc/research-api-specs-query-videos. While this indicator is likely confounded by VPN usage and other factors, it represents our best method for assessing a post’s country of origin. As detailed later in Section 4.1, our method has slight variation in regional representativeness, so we adjust counts based on known statistics from our method. We explain precisely how we do this in Section 4.3.
@@ 2.3 Global usage
Figure 4 shows a regression plot of country population against post count in our dataset ($R^{2}=0.69$, $p=0.00$). We see a strong fit to the data, evidence that we have successfully collected a globally representative dataset. We see that China, India, and Cuba, have the least posts in our dataset per capita, based on the residual of a fitted linear regression. We note that while TikTok was previously extremely popular in India [77], there are now very few posts originating from India (as deemed by TikTok), likely due to the nationwide censorship policy on Chinese applications there [118]. Of course, there could still be posts produced from India via virtual private networks (VPNs), which would be tagged with a different origin country by TikTok. TikTok is not officially available in China [83], and Cuba has heavy internet restrictions [67]. Monaco, Qatar, and the UAE, have the most posts in our dataset per capita.
The post metadata also contains an attribute specifying whether the post media contains artificial-intelligence (AI)-generated content, whether tagged by the user or by the platform, including AI-generated images, video, audio, and filters  [^4] https://support.tiktok.com/en/using-tiktok/creating-videos/ai-generated-content. In our 24-hour dataset, the percentage of content that is AI-generated according to TikTok is 0.5%. This percentage varies by country (Fig. 5). We find evidence of other posts containing AI-generated that are not tagged as such by TikTok, indicating the true statistic may be higher. However, we will leave a more thorough investigation of this important topic for future work.
Figure 5: Share of posts tagged as AI generated content. Areas in grey are countries where we had less than 50 posts available.
Using the locationCreated attribute, we can also infer the distribution of when TikTok posts are created in the local time, see Figure 6. In the case of countries spanning multiple time-zones, we use the mean hour from each time-zone associated with the country from the Olson time-zone database  [^5] https://github.com/stub42/pytz/blob/master/tz/zone.tab. Inaccuracies are inevitable due to lack of sub-country location information, and countries spanning multiple time-zones, but the general trends should be preserved nonetheless.
Figure 6: Distribution of local time posting frequency, found by looking at post time combined with the locationCreated tag.
@@ 2.4 Video Media
Figure 7: View of the topics covered by the posts, with each small dot representing a video, and distance between dots representing video content similarity. Common topics of videos are coloured and labelled with a description describing the topic.
@@
The topic analysis completed here is not meant to be comprehensive. But by releasing the unsupervised topics discovered, we provide researchers an entry-point to contemplating the study and engagement of particular domains within TikTok.
@@ 2.4.1 Presence of Children
The use of TikTok by children has caused concern for safety on the platform [119], but without knowing what percentage of content on the platform features children, we cannot know the true magnitude of the issue, or make connections between localised policy and content outcomes.
@@
We trained a classifier to detect if a video contains a child, with a mean F1 of 0.86 from 5-fold cross-validation on our train/test-set (precision 0.88, recall 0.81). We provide data, model selection, and training details in Section 4.4.2. We used this classifier to find that 18.00% of the videos in our 24 hour dataset contain children (95% confidence interval: [13.08%, 24.62%] from bootstrapped confidence intervals, see Section 4.4.2 for details). This is a strikingly high statistic, and underscores how deeply embedded children are in platform activity. To the best of our knowledge, no comparable large-scale studies exist for TikTok or other platforms that would allow direct benchmarking or comparison of this figure.
@@
Figure 8: Choropleth of share of videos from our 1 minute per hour dataset containing one or more children, divided by the share of the population under the age of 14 [106] (to avoid showing simply a map of countries with a youthful population), plotted by country. Areas in grey are countries where we had less than 50 posts or videos available.
@@
Figure 9: Scatter plot of countries, for each country showing the share of the population under the age of 14 against the share of videos from that country containing one or more children, as determined by our video classifier. We have removed countries where there was insufficient data simply by removing countries with below 50 posts in the dataset. We label the top 2 and bottom 2 countries with the highest residual from the linear regression, effectively showing countries with the most overrepresented and under-represented children in TikTok posts, respectively. Red points correspond to the country name closest to them.
@@
We observe variability in the fraction of videos containing children by country. Figure 8 shows a choropleth of share of videos containing one or more children per country, adjusted by the share of the population under the age of 14  [^6] https://population.un.org/wpp/ (to avoid showing simply a map of countries with a youthful population). We removed countries where there were less than 50 posts in the dataset. This effectively shows countries where there is a culture of posting videos containing children. Figure 9 shows a scatter plot of countries comparing the share of the population under 14, and the share of videos containing one or more children - a linear trend is evident. We label the top 2 and bottom 2 countries with the highest residual from the linear regression ($R^{2}=0.51$, $p=0.00$), effectively showing countries with the most overrepresented and under-represented children in TikTok posts, respectively. The variation we observe may reflect broader differences in cultural norms around public sharing of children’s images, the popularity of family-centred content, differences in platform access and usage patterns, and differences in the implementation of child protection policies across countries.